# Wasserstein Distributionally Robust Shallow Convex Neural Networks

### v2

@@ 3 Wasserstein Distributionally Robust Shallow Convex Neural Networks
[…]
Now, recall that $\mathcal{S}$ is known a priori and can be generated randomly to approximate the hyperplane arrangement of ^$\hat{\mathbb{P}}_{N}$. As such, $\mathcal{S}$ do not bring additional information to $\mathbf{x}_{j}$; indeed $\hat{\mathbf{x}}_{j}$ is purely an expansion of $\mathbf{x}_{j}$ with a possible $\operatorname{ReLU}$ activation pattern. Effectively, we map the original samples to a higher dimension and apply the Wasserstein distributionally robust regularization on this higher representation which is closely related to the original feature space.
We formulate the non-tractable Wasserstein distributionally robust training problem of the SCNN by centering the Wasserstein ball on $\hat{\mathbb{F}}_{N}^{\mathcal{S}}$:
$\inf_{\bm{\beta}}\sup_{\mathbb{Q}\in\Omega}$ $\mathbb{E}^{\mathbb{Q}}[g_{\bm{\beta}}(\hat{\mathbf{z}})|\mathcal{S}]$ (WDR-SCNNT)
@@
$\Omega=\{\mathbb{Q}\in\mathcal{P}(\hat{\mathcal{Z}}):W_{\|\cdot\|,t}(\mathbb{Q},\hat{\mathbb{F}}_{N}^{\mathcal{S}})\leq\epsilon\},$
where $\hat{\mathbf{z}}\in\hat{\mathcal{Z}}$ and $\mathcal{P}(\hat{\mathcal{Z}})$ is the space of all possible distributions with support on $\hat{\mathcal{Z}}=\mathbb{R}^{2Pd+1}$. Note that $g_{\bm{\beta}}(\hat{\mathbf{z}})=\left|\bm{\beta}^{\top}\left(\operatorname{vec}(\mathbf{x}_{j}\mathbf{d}_{j}^{\top}),\operatorname{vec}(\mathbf{x}_{j}\mathbf{d}_{j}^{\top}),y_{j}\right)\right|$ as defined in Lemma 2.
@@
    $\min_{\bm{\nu},\bm{\omega},a,\mathbf{c}}\quad$ $\epsilon a$ $+\frac{1}{N}\sum_{j=1}^{N}c_{j}$ (WDR1-SCNNT)
    s.t. ${\bm{\beta}}_{k}$ $\leq a\quad$ $\forall k\in\llbracket 2Pd+1\rrbracket$
    $-{\bm{\beta}}_{k}$ $\leq a\quad$ $\forall k\in\llbracket 2Pd+1\rrbracket$
@@
    ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bm{\beta}}}$ ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}=\left(\operatorname{vec}(\bm{\nu}),\operatorname{vec}(-\bm{\omega}),-1\right)}}$
@@
    ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bm{\beta}}}$ ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}=\left(\operatorname{vec}(\bm{\nu}),\operatorname{vec}(-\bm{\omega}),-1\right)}}$
@@
Proof: Lemma 1 states that the training problem, with respect to the modified samples $\hat{\mathbf{z}}$, is similar to a constrained linear regression over a basis expansion of the original feature space. This problem is convex in its objective and its constraints. Lemma 2 shows that the distribution of the modified samples is i.i.d. when conditioned on $\mathcal{S}$. Because (WDR-SCNNT) centers the Wasserstein ball on $\hat{\mathbb{F}}_{N}^{\mathcal{S}}$, by invoking Lemmas 1 and 2 regarding the modified sample space, we can reexpress (WDR-SCNNT) to the tractable DRO formulation (TWDRO). The final forms then follow from Theorem 6.3 of  Mohajerin Esfahani and Kuhn (2018) and are inherited from WaDiRo linear regressions (Chen and Paschalidis 2020, Chapter 4). The detailed proof, adapted to shallow convex neural networks, is provided in Appendix A for completeness. $\square$
We remark that (WDR1-SCNNT) and (WDR2-SCNNT) can be easily modified to take into account the bias weights of the hidden and output layers of the SNN. The process is as follows: (i) extend $\mathbf{X}$ with a column of ones which is equivalent to adding bias weights in the hidden layer; (ii) let $\bm{\beta}$ be defined as $\left(\operatorname{vec}(\bm{\nu}),\operatorname{vec}(-\bm{\omega}),b,-1\right)$ and $\hat{\mathbf{z}}_{j}$ as $\left(\operatorname{vec}(\mathbf{x}_{j}\mathbf{d}_{j}^{\top}),\operatorname{vec}(\mathbf{x}_{j}\mathbf{d}_{j}^{\top}),1,y_{j}\right)\ \forall j\in\llbracket N\rrbracket$ where $b\in\mathbb{R}$ is the bias of the sole output neuron. We also remark that both formulations of Proposition 1 are fairly fast to solve with open-source convex solvers, e.g., Clarabel (Goulart and Chen 2024), if the maximal number of neurons dictated by $\tilde{\mathcal{D}}_{X}$ and the size of the training set are reasonable in regards to the available computational power, e.g., in our case less than 300 neurons and 3000 samples were required. We provide time comparisons in Section B.3 and identify acceleration methods from the literature that could be adapted to our formulation in Section 9. Moreover, with only two hyperparameters, i.e., the maximal number of neurons and the Wasserstein ball radius, the hyperparameter search space is significantly reduced compared to standard non-convex training programs. As such, they contribute to a lighter MLOps pipeline by diminishing the computational effort required during hyperparameter tuning. Recall that the weights from the standard non-convex SNN training problem can be retrieved via (2) and (3) even though they are not required to obtain accurate predictions as the SCNN representation is sufficient.
@@ 4 Out-of-sample performance
The norm of modified samples from $\mathbb{F}^{\mathcal{S}}$ are bounded above: $\|{\hat{\mathbf{z}}}\|\leq\hat{R}<+\infty$.
@@
The dual norm of the weights is bounded above: $\sup_{\bm{\beta}}\|\bm{\beta}\|_{*}={B}_{*}<+\infty$.
These assumptions are reasonable as we aim for real-world applications which, by nature, generally satisfy them.
@@
Under Assumptions 1, 2, and 3, and considering an $\ell_{1}$-loss function, with probability at least $1-\delta$ with respect to the sampling, the expected out-of-sample error of WaDiRo-SCNNs on $\mathbb{F}^{\mathcal{S}}$ is:
$\mathbb{E}^{\mathbb{F}^{\mathcal{S}}}\left[\left|{\bm{\beta}}^{\star\top}{\hat{\mathbf{z}}}\right|\middle|\mathcal{S}\right]\leq\frac{1}{N}\sum_{j=1}^{N}\left|{\bm{\beta}}^{\star\top}{\hat{\mathbf{z}}}_{j}\right|+\frac{2{B}_{*}\hat{R}}{\sqrt{N}}+B_{*}\hat{R}\sqrt{\frac{8\ln(2/\delta)}{N}},$
for a modified sample $\hat{\mathbf{z}}$ from the true distribution, that was not available for training, where ${\bm{\beta}}^{\star}$ are trained weights obtained from either (WDR1-SCNNT) or (WDR2-SCNNT).
@@
Proof: By Lemma 1, the WaDiRo-SCNN’s training loss function with respect to $\hat{\mathbb{F}}^{\mathcal{S}}_{N}$ belongs to the following family:
=$\mathcal{H}=\left\{(\mathbf{x},y)\to h_{\bm{\beta}}(\mathbf{x},y):h_{\bm{\beta}}(\mathbf{x},y)=\left|y-\mathbf{x}^{\top}\bm{\beta}\right|\right\}.$
@@
Interestingly, we can obtain similar performance bounds with respect to the original distribution $\mathbb{P}$ instead of the modified one. We now make further assumptions.
@@
Consider $\psi_{1}=B_{*}2PR$ for (WDR1-SCNNT) and $\psi_{2}=B_{*}\sqrt{2P}R$ for (WDR2-SCNNT), upper bounds on the prediction error of the WaDiRo-SCNN for any sample from the sample space. Then, we can rewrite the out-of-sample performance of trained WaDiRo-SCNNs as follows.
@@
Under Assumptions 1, 3, and 4, and considering an $\ell_{1}$-loss function, with probability at least $1-\delta$ with respect to the sampling, the expected out-of-sample error of WaDiRo-SCNNs on $\mathbb{P}^{\mathcal{S}}$ is:
$\mathbb{E}^{\mathbb{P}^{\mathcal{S}}}\left[\left|{\bm{\beta}}^{\star\top}{\hat{\mathbf{z}}}\right|\middle|\mathcal{S}\right]\leq\frac{1}{N}\sum_{j=1}^{N}\left|{\bm{\beta}}^{\star\top}{\hat{\mathbf{z}}}_{j}\right|+\frac{2\psi_{l}}{\sqrt{N}}+B_{*}\hat{R}\sqrt{\frac{8\ln(2/\delta)}{N}},$ (5)
for a modified sample $\hat{\mathbf{z}}$ from the true distribution given $\mathcal{S}$, that was not available for training, where ${\bm{\beta}}^{\star}$ are the optimal weights obtained from either (WDR1-SCNNT) or (WDR2-SCNNT), $\mathbb{P}^{\mathcal{S}}$ is the conditioned underlying distribution, and where $l\in\{1,2\}$.
@@
$|\bm{\beta}^{\top}\hat{\mathbf{z}}|$ $\leq\|\bm{\beta}\|_{2}\|\hat{\mathbf{z}}\|_{2}$ (Hölder’s inequality)
$\leq B_{*}\sqrt{\sum_{i=1}^{P}\sum_{k=1}^{d}2(\mathbf{x}_{k}\text{1}(\mathbf{x}^{\top}\mathbf{s}_{i}\geq 0))^{2}+y^{2}}$
$\leq B_{*}\sqrt{\sum_{i=1}^{P}\sum_{k=1}^{d}2(\mathbf{x}_{k})^{2}+y^{2}}$ (Full activation)
@@
$|\bm{\beta}^{\top}\hat{\mathbf{z}}|$ $\leq\|\bm{\beta}\|_{\infty}\|\hat{\mathbf{z}}\|_{1}$ (Hölder’s inequality)
$\leq B_{*}\sum_{i=1}^{P}\sum_{k=1}^{d}2|(\mathbf{x}_{k}\text{1}(\mathbf{x}^{\top}\mathbf{s}_{i}\geq 0))|+|y|$
@@
Thus, by adapting Theorem 8 of Bartlett and Mendelson (2002) to loss functions mapping between $[0,B_{*}\hat{R}]$ and using the fact that $\mathcal{R}_{N}(\mathcal{G})$ is greater than the Rademacher complexity term defined in Theorem 8, we obtain the expression from (5). $\square$
We remark that the expected out-of-sample error reduces to the training error as the number of training samples grows. We also remark that the tighter the bounds are on the sample space, i.e., $\hat{R},\ R,\ S,\text{ and}\ T$, the tighter the guarantees are. Proof of Proposition 3 also leads to a probabilistic guarantee similar to the one in Proposition 2.
@@ 5 Physics-constrained shallow convex neural networks
$\bm{\nu}_{i},\bm{\omega}_{i}\in\mathcal{K}_{i}$ $\forall i\in\llbracket P\rrbracket,$
@@
where in (6), (7), and (8), with respect to sample $j$ and $\forall i\in\llbracket j\rrbracket$, $u_{j}\in\mathbb{R}$ is an upper bound, $l_{j}\in\mathbb{R}$ is a lower bound, and $r_{j}\in\mathbb{R}$ is a ramping bound. Additionally, (8) is a bounded first-order relation between input and output history with $\underline{\xi}_{j}\in\mathbb{R},\bar{\xi}_{j}\in\mathbb{R},\bm{\eta}_{i}\in\mathbb{R}^{d},\gamma_{i}\in\mathbb{R}$, e.g., a discrete-time dynamical equation. These examples are linear or easily linearizable but we stress that any convex constraint can be included. In the case of non-differentiable convex functions, the numerical solver must be chosen accordingly or a subgradient method be implemented Boyd and Park (2014). We remark that the WaDiRo-SCNN training problem can be similarly modified to consider physical constraints by adding $\hat{y}_{j}\in\mathcal{Y}_{j}\ \forall j\in\llbracket N\rrbracket$ to the constraints of (WDR1-SCNNT) and (WDR2-SCNNT).
@@
from which we find (PCSCNNT). $\square$
We remark that a similar approach can be used for the regularized problem.

### v3

@@ 3 Wasserstein Distributionally Robust Shallow Convex Neural Networks
$\hat{\mathbf{z}}_{j}$ $=(\hat{\mathbf{x}}_{j},y_{j})=\left(\operatorname{vec}(\mathbf{x}_{j}\mathbf{d}_{j}^{\top}),\operatorname{vec}(\mathbf{x}_{j}\mathbf{d}_{j}^{\top}),y_{j}\right)=\left((\mathbf{x}_{j},\ \mathbf{0}_{d},\ \mathbf{x}_{j}),(\mathbf{x}_{j},\ \mathbf{0}_{d},\ \mathbf{x}_{j}),y_{j}\right).$
Now, recall that $\mathcal{S}$ is known a priori and can be generated randomly to approximate the hyperplane arrangement of $\hat{\text{P}}_{N}$. As such, $\mathcal{S}$ do not bring additional information to $\mathbf{x}_{j}$; indeed $\hat{\mathbf{x}}_{j}$ is purely an expansion of $\mathbf{x}_{j}$ with a possible $\operatorname{ReLU}$ activation pattern. Effectively, we map the original samples to a higher dimension and apply the Wasserstein distributionally robust regularization on this higher representation, which is closely related to the original feature space. $\square$
We formulate the intractable Wasserstein distributionally robust training problem of the SCNN by centering the Wasserstein ball on $\hat{\text{F}}_{N}^{\mathcal{S}}$:
$\inf_{\bm{\beta}}\sup_{\text{Q}\in\Omega}\$$\mathbb{E}^{\text{Q}}[g_{\bm{\beta}}(\hat{\mathbf{z}})|\mathcal{S}]$ (WDR-SCNNT)
@@
[…]
$\Omega=\{\text{Q}\in\mathcal{P}(\hat{\mathcal{Z}}):W_{\|\cdot\|,t}(\text{Q},\hat{\text{F}}_{N}^{\mathcal{S}})\leq\epsilon\}.$
Note that $g_{\bm{\beta}}(\hat{\mathbf{z}})=\left|\bm{\beta}^{\top}\left(\operatorname{vec}(\mathbf{x}\mathbf{d}^{\top}),\operatorname{vec}(\mathbf{x}\mathbf{d}^{\top}),y\right)\right|$ as defined in the proof of Lemma 2.
@@
    $\min_{\bm{\nu},\bm{\omega},a,\mathbf{c}}\quad$ $\epsilon a\$$+\frac{1}{N}\sum_{j=1}^{N}c_{j}$ (WDR1-SCNNT)
    s.t. ${{\beta}}_{k}$ $\leq a\quad$ $\forall k\in\llbracket 2Pd+1\rrbracket$
    $-{{\beta}}_{k}$ $\leq a\quad$ $\forall k\in\llbracket 2Pd+1\rrbracket$
@@
[…]
@@
[…]
@@
Proof: Lemma 1 states that the training problem, with respect to the modified samples $\hat{\mathbf{z}}$, is similar to a constrained linear regression over a basis expansion of the original feature space. This problem is convex in its objective and its constraints. Lemma 2 shows that the distribution of the modified samples is i.i.d. when conditioned on $\mathcal{S}$. Because (WDR-SCNNT) centres the Wasserstein ball on $\hat{\text{F}}_{N}^{\mathcal{S}}$, by invoking Lemmas 1 and 2 regarding the modified sample space, we can reexpress (WDR-SCNNT) to the tractable DRO formulation (TWDRO). The final forms then follow from Theorem 6.3 of  Mohajerin Esfahani and Kuhn (2018) and are inherited from WaDiRo linear regressions (Chen and Paschalidis 2020, Chapter 4). The detailed proof, adapted to shallow convex neural networks, is provided in Appendix A for completeness. $\square$
We remark that (WDR1-SCNNT) and (WDR2-SCNNT) can be easily modified to take into account the bias weights of the hidden and output layers of the SNN. The process is as follows: (i) extend $\mathbf{X}$ with a column of ones which is equivalent to adding bias weights in the hidden layer; (ii) let $\bm{\beta}$ be defined as $\left(\operatorname{vec}(\bm{\nu}),\operatorname{vec}(-\bm{\omega}),b,-1\right)$ and $\hat{\mathbf{z}}_{j}$ as $\left(\operatorname{vec}(\mathbf{x}_{j}\mathbf{d}_{j}^{\top}),\operatorname{vec}(\mathbf{x}_{j}\mathbf{d}_{j}^{\top}),1,y_{j}\right)\ \forall j\in\llbracket N\rrbracket$ where $b\in\mathbb{R}$ is the bias of the sole output neuron. We also remark that both formulations of Proposition 1 are fairly fast to solve with open-source convex solvers, e.g., Clarabel (Goulart and Chen 2024), if the maximal number of neurons dictated by $\tilde{\mathcal{D}}_{\mathbf{X}}$ and the size of the training set are reasonable in regards to the available computational power, e.g., in our case less than 300 neurons and 3000 samples were required. We provide time comparisons in Section B.3 and identify acceleration methods from the literature that could be adapted to our formulation in Section 9. Moreover, with only two hyperparameters, i.e., the maximal number of neurons and the Wasserstein ball radius, the hyperparameter search space is significantly reduced compared to standard non-convex training programs. As such, they contribute to a lighter MLOps pipeline by diminishing the computational effort required during hyperparameter tuning. Recall that the weights from the standard non-convex SNN training problem can be retrieved via (2) and (3) even though they are not required to obtain accurate predictions as the SCNN representation is sufficient.
@@ 4 Out-of-sample performance
The norm of modified samples from $\text{F}^{\mathcal{S}}$ are bounded above: $\|{\hat{\mathbf{z}}}\|\leq\hat{R}$$\infty$.
@@
The dual norm of the trained weights, $\bm{\beta}^{\star}$, is bounded above: $\|\bm{\beta}^{\star}\|_{*}\leq{B}_{*}<+\infty$.
These assumptions are reasonable as we aim for real-world applications which, by nature, generally satisfy them, and trained weights are finite.
@@
Under Assumptions 1, 2, and 3, and considering an $\ell_{1}$-loss function, with probability at least $1-\delta$ with respect to the sampling, the expected out-of-sample error of WaDiRo-SCNNs on $\text{F}^{\mathcal{S}}$ is:
$\mathbb{E}^{\text{F}^{\mathcal{S}}}\left[\left|{\bm{\beta}}^{\star\top}{\hat{\mathbf{z}}}\right|\middle|\mathcal{S}\right]\leq\frac{1}{N}\sum_{j=1}^{N}\left|{\bm{\beta}}^{\star\top}{\hat{\mathbf{z}}}_{j}\right|+\frac{2{B}_{*}\hat{R}}{\sqrt{N}}+B_{*}\hat{R}\sqrt{\frac{8\ln(2/\delta)}{N}},$
for a modified sample $\hat{\mathbf{z}}$ from the true distribution, that was not available for training, where $\bm{\beta}$$^{\star}$ are trained weights obtained from either (WDR1-SCNNT)or (WDR2-SCNNT)).
@@
Proof: By Lemma 1, the WaDiRo-SCNN’s training loss function with respect to $\hat{\text{F}}^{\mathcal{S}}_{N}$ belongs to the following family:
$\mathcal{H}=\left\{(\mathbf{x},y)\to h_{\bm{\beta}}(\mathbf{x},y):h_{\bm{\beta}}(\mathbf{x},y)=\left|y-\mathbf{x}^{\top}\bm{\beta}\right|\right\}.$
@@
Interestingly, we can obtain similar performance bounds with respect to the original distribution P instead of the modified one. We now make further assumptions.
@@
Consider $\psi_{1}=B_{*}2PR$ for (WDR1-SCNNT) and $\psi_{2}=B_{*}\sqrt{2P}R$ for (WDR2-SCNNT), upper bounds on the prediction error of the WaDiRo-SCNN for any sample from the sample space. Then, we can rewrite the out-of-sample performance of trained WaDiRo-SCNNs with respect to $\text{P}^{\mathcal{S}}$, the conditioned underlying distribution, as follows.
@@
Under Assumptions 1, 3, and 4, and considering an $\ell_{1}$-loss function, with probability at least $1-\delta$ with respect to the sampling, the expected out-of-sample error of WaDiRo-SCNNs on $\text{P}^{\mathcal{S}}$ is:
$\mathbb{E}^{\text{P}^{\mathcal{S}}}\left[\left|{\bm{\beta}}^{\star\top}{\hat{\mathbf{z}}}\right|\middle|\mathcal{S}\right]\leq\frac{1}{N}\sum_{j=1}^{N}\left|{\bm{\beta}}^{\star\top}{\hat{\mathbf{z}}}_{j}\right|+\frac{2\psi_{l}}{\sqrt{N}}+B_{*}\hat{R}\sqrt{\frac{8\ln(2/\delta)}{N}},$ (5)
for a modified sample $\hat{\mathbf{z}}$ from the true distribution given $\mathcal{S}$, that was not available for training, where $\bm{\beta}$$^{\star}$ are the optimal weights obtained from either (WDR1-SCNNT)) or (WDR2-SCNNT)) and where $l\in\{1,2\}$.
@@
[…]
$\leq B_{*}\sqrt{\sum_{i=1}^{P}\sum_{k=1}^{d}2({x}_{k}\text{1}(\mathbf{x}^{\top}\mathbf{s}_{i}\geq 0))^{2}+y^{2}}$
$\leq B_{*}\sqrt{\sum_{i=1}^{P}\sum_{k=1}^{d}2({x}_{k})^{2}+y^{2}}$ (Full activation)
@@
[…]
$\leq B_{*}\sum_{i=1}^{P}\sum_{k=1}^{d}2|({x}_{k}\text{1}(\mathbf{x}^{\top}\mathbf{s}_{i}\geq 0))|+|y|$
@@
Thus, by adapting Theorem 8 of Bartlett and Mendelson (2002) to loss functions mapping between $0,B_{*}\hat{R}$] and using the fact that $\mathcal{R}_{N}(\mathcal{G})$ is greater than the Rademacher complexity term defined in Theorem 8, we obtain the expression from (5). $\square$
We remark that the expected out-of-sample error reduces to the training error as the number of training samples grows. We also remark that the tighter the bounds are on the sample space, i.e., $\hat{R},\ R,\ S,\text{ and}\ T$, the tighter the guarantees are. The proof of Proposition 3 also leads to a probabilistic guarantee similar to the one in Proposition 2.
@@ 5 Physics-constrained shallow convex neural networks
$\bm{\nu}_{i},\bm{\omega}_{i}\in\mathcal{K}_{i}\$$\forall i\in\llbracket P\rrbracket,$
@@
where in (6), (7), and (8), with respect to sample $j$ and $\forall i\in\llbracket j\rrbracket$, $u_{j}\in\mathbb{R}$ is an upper bound, $l_{j}\in\mathbb{R}$ is a lower bound, and $r_{j}\in\mathbb{R}$ is a ramping bound. Additionally, (8) is a bounded first-order relation between input and output history with $\underline{\xi}_{j}\in\mathbb{R},\bar{\xi}_{j}\in\mathbb{R},\bm{\eta}_{i}\in\mathbb{R}^{d},\gamma_{i}\in\mathbb{R}$, e.g., a discrete-time dynamical equation or PDE. These examples are linear or easily linearizable but we stress that any convex constraint can be included. In the case of non-differentiable convex functions, the numerical solver must be chosen accordingly or a subgradient method be implemented Boyd and Park (2014). We remark that the WaDiRo-SCNN training problem can be similarly modified to consider physical constraints by adding $\hat{y}_{j}\in\mathcal{Y}_{j}\ \forall j\in\llbracket N\rrbracket$ to the constraints of (WDR1-SCNNT) and (WDR2-SCNNT).
@@
from which we find (PCSCNNT). We remark that a similar approach can be used for the regularized problem and that the converse of the proposition can be obtained following the same idea. $\square$